OpenStudy (anonymous):

find the distance between P(-2,1) and the line with equation x-2y+4=0

OpenStudy (anonymous):

its zero

OpenStudy (anonymous):

yes and do you want me to show you how?

OpenStudy (anonymous):

k so first you want to plug in the x and y p(-2,1) ( x,y ) x-2y+4=0 -2-2(1)+4=0 -2-2+4=0 -4+4=0 0=0